4th and 7th term of an AP are 9 and 21 respectively. Find the 1st eight terms of the sequence
Answer by richwmiller(17219)

an = a + (n - 1)d
21= a + 6d
9 = a + 12
12=3d
d=4
a=-3
-3+4=1 etc
-3,1,5,9,13,17,21,25